OpenStudy (anonymous):

Solve for x. |2x + 2| = 14

OpenStudy (ivettef365):

since it is an absolute value then you have two answers: 2x +2 = 14 and 2x +2 = =14

OpenStudy (ivettef365):

I meant 2x + 2 = -14

OpenStudy (ivettef365):

solve for x

OpenStudy (anonymous):

6 and 8 or 8 and 6

OpenStudy (ivettef365):

2x +2 = -14 -2 - 2 ------------ 2x = -16 so x = -8 so 6 and -8

OpenStudy (ivettef365):

what that means is that the equation {2x+2} = 14 is true when x = 6 or -8, plug the numbers to see if is true

OpenStudy (ivettef365):

so you have {2(6)+2} = 14 {12+2} = 14 {14} = 14 TRUE

OpenStudy (ivettef365):

now the other {2(-8)+2} = 14 {-16+2} = 14 {-14} = 14 TRUE